Define a metric space .Give an example of a metric space


1
Expert's answer
2021-08-03T06:51:44-0400

A metric space is an ordered pair "{\displaystyle (M,d)}",  where "{\displaystyle M}" is a set and "{\displaystyle d}"  is a metric on "{\displaystyle M}", i.e., a function

"{\displaystyle d\,\colon M\times M\to \mathbb {R} }" such that for any "{\displaystyle x,y,z\in M}" , the following holds:

  1. "{\displaystyle d(x,y)=0\iff x=y}" (identity of indiscernibles)
  2. "{\displaystyle d(x,y)=d(y,x)}" (symmetry)
  3. "{\displaystyle d(x,z)\leq d(x,y)+d(y,z)}" (subadditivity or triangle inequality).

An example of a metric space is "(\mathbb R,d)," where "d(x,y)=|x-y|." Indeed,

  1. "{\displaystyle d(x,y)=0\iff |x-y|=0 \iff x-y=0\iff x=y}"
  2. "{\displaystyle d(x,y)=|x-y|=|-(y-x)|=|y-x|=d(y,x)}"
  3. "{\displaystyle d(x,z)=|x-z=|x-y+y-z|\leq|x-y|+|y-z|= d(x,y)+d(y,z)}"